# Fine-Tuning Image-Conditional Diffusion Models is Easier than You Think

### v1

@@ 2 Related Work
Monocular Depth Estimation.
Monocular depth estimation involves predicting a pixel-wise depth map of a scene from a single image. The simplest depth representation is relative depth, which is purely ordinal, providing only a depth-wise ranking of objects in the scene [7, 8]. In contrast, the most comprehensive representation is metric depth, which requires modeling the focal length to account for different cameras, introducing additional uncertainty [4, 36, 54, 23].
A middle ground between relative and metric depth is affine-invariant depth, which is equivalent to metric depth up to an unknown global scale and shift of the scene. This is the representation of choice for a wide range of monocular depth estimation methods [56, 39, 38, 12, 24, 51, 52]. Unlike metric depth, affine-invariant depth is independent of the focal length and is easier to regress in unfamiliar scenes, where no object can serve as a metric reference. However, it still preserves the distance ratios between objects. Metric depth can be recovered from affine-invariant depth by anchoring it with sparse, known depth values or by explicitly estimating the missing scale and shift [55].
@@
The transition from CNNs to ViTs [10] in DPT [38] and Omnidata [12] marked another key advancement in the field. Recent methods such as Depth Anything [51, 52] and Metric3D [54, 23] have followed the success of MiDaS by utilizing the high-capacity ViT-g DINOv2 [35] backbone and training on vast datasets—62M and 16M samples, respectively—one to two orders of magnitude larger than previous work. Notably, Depth Anything retains a simple DPT architecture, but combining a DINOv2 initialization and a large training dataset enables it to generalizate to in-the-wild scenarios well. Metric3D additionally utilizes the focal length to boost performance, which however limits its training data to scenes with known focal length.
Monocular Normal Estimation.
Surface normal estimation involves predicting the orientation of surfaces in a scene from an image, resulting in a 3D vector representing the surface’s orientation for each pixel. Hoiem et al. [21, 22] were among the pioneers to introduce learning-based approaches for surface normal estimation. Since then, deep learning methods have become dominant, with many notable contributions [49, 13, 3, 1, 12, 24, 2]. Among these, Omnidata [12] stands out for training a UNet-based model on a large-scale dataset of 12M images captured in diverse environments. Its successor, Omnidata v2 [24], advanced the field by transitioning to a ViT-based architecture with a DPT head, incorporating sophisticated 3D-aware data augmentations to enhance generalization. In contrast, DSINE [2] adopts a more data-efficient approach that focuses on introducing new inductive biases to enhance performance. Lastly, Metric3D v2 [23], mentioned earlier, is also capable of predicting surface normals in addition to depth.
Diffusion Models for Geometry Estimation.
Several recent generative text-to-3D methods [32, 37] explicitly produce multi-view depth and normal maps. However, these methods focus on isolated single-object scenarios, making them unsuitable for complex, in-the-wild environments.
Other approaches have utilized diffusion models for scene-level depth estimation [44, 43, 11]. Among these, VPD [57] leverages Stable Diffusion [41] as both an image and a text feature extractor, incorporating a depth regression head that utilizes multi-scale image feature maps alongside text-to-image cross-attention maps. However, these models have not demonstrated robust generalization.
@@
In our work, we observe that Marigold and follow-up methods, aside from DepthFM and Marigold LCM, which are designed for few-step prediction, suffer from a flawed implementation [30] of the DDIM [47] inference pipeline that prevents them from functioning effectively in the few-step regime. Furthermore, although the denoising diffusion fine-tuning objective used by Marigold and follow-up works for depth and normals estimation has shown effectiveness, we find it to be neither a key factor nor particularly advantageous compared to task-specific end-to-end fine-tuning.
@@ 3.2 Marigold
Marigold casts depth estimation as a conditional latent diffusion process, allowing it to build upon large pretrained diffusion models such as Stable Diffusion [26]. Marigold is conditioned on images, so following the argument of Sec. 3.1, we expect its single-step prediction to be a sensible, but blurry depth map.
Training.
Fig. 3 shows an overview of Marigold’s training procedure. Marigold adapts the SD v2 [41] UNet architecture for conditional depth map generation, using $\bm{\mathbf{v}}$-parametrization [42] during training. The training objective is formulated in the latent space of the frozen SD VAE. The GT depth map $\bm{\mathbf{d}}^{*}$ is replicated along the channel dimension to conform to the 3-channel inputs of the VAE and encoded as $\bm{\mathbf{z}}^{*}=\mathcal{E}(\bm{\mathbf{d}}^{*})$. Similarly, the RGB latent is $\bm{\mathbf{x}}=\mathcal{E}(\bm{\mathbf{I}}_{\text{RGB}})$. During training, noise is added only to the depth latent to get $\bm{\mathbf{z}}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{\mathbf{z}}^{*}+\sqrt{1-\bar{\alpha}_{t}}\bm{\mathbf{\epsilon}}$. The UNet receives the concatenated latents and the timestep $t$ as inputs and predicts $\bm{\mathbf{\hat{v}}}_{t}=\bm{\mathbf{\hat{v}}}_{\theta}([\bm{\mathbf{z}}_{t},\bm{\mathbf{x}}],t)$. The first convolutional layer is duplicated to accomodate the larger number of input channels [26].
@@
Fixing Single-Step Inference.
We now turn to analyze the behavior of Marigold during inference. First, we show the pixel-wise standard deviation across steps for Marigold’s default 50-step inference in  Fig. 2. We observe very small differences for almost all pixels, indicating that the model changes predictions very little during inference; in particular, this means that the first prediction of the 50-step schedule already corresponds closely to the final output. While we would expect that the single-step output should be similar to the first step of the 50-step schedule, as shown in Fig. 2 it rather corresponds to pure noise.
@@
Fixing the flaw is simple: we need to align the timestep with the noise level. To do this, we can use the trailing setting as proposed in recent work [30] for image-generative models. However, we emphasize that while this setting provided only slight improvements for image generation [30], this setting is crucial for single-step inference in models such as Marigold. In Fig. 2 and Fig. 2, we compare the outputs of the same model, using the flawed and the fixed inference schedule, respectively. Clearly, the fixed inference process produces sensible predictions, while the original does not.
@@ 5 Experimental Setup
Training Datasets.
To allow for direct comparison with Marigold [26], we use the same training datasets: Hypersim [40], which consists of photorealistic indoor scenes, and Virtual KITTI 2 [6], which covers driving scenarios. Both datasets are fully synthetic and provide high-quality ground-truth annotations.
@@
Evaluation Datasets.
We evaluate the fine-tuned models on commonly used benchmarks for monocular depth estimation. NYUv2 [46] and ScanNet [9] provide RGB-D data of indoor environments captured with Kinect cameras. ETH3D [45] and DIODE [48] consist of both indoor and outdoor scenes, derived from LiDAR sensors. KITTI [17] contains outdoor driving scenes captured by vehicle-mounted cameras and LiDAR sensors. For surface normal estimation, we evaluate on NYUv2, ScanNet, and additionally on iBims-1 [28], a high-quality indoor RGB-D dataset, as well as Sintel [5], a synthetic outdoor dataset.
Evaluation Protocol.
All evaluations are conducted in the zero-shot setting. We evaluate affine-invariant depth predictions using the standard approach, which involves the same scale and shift optimization between the predicted depth and the ground truth as in the loss computation [39]. Following established conventions, we report the mean absolute relative error (AbsRel), defined as the average relative difference between the ground-truth depth and the aligned predicted depth at each pixel, as well as the $\delta$1 accuracy, which is the percentage of pixels where the ratio of the aligned predicted depth to the ground truth (and its inverse) is less than 1.25.
@@
Implementation Details.
For depth estimation, we use the official Marigold checkpoint, whereas for normal estimation, we train a model with the same training setup as Marigold’s depth estimation, encoding normal maps as 3D vectors in the color channels.
Unless noted otherwise, we follow Marigold’s hyperparameters. While a more extensive hyperparameter search might further improve results, the Marigold defaults performed well in our experiments. We train all models for 20K iterations using the AdamW optimizer [33] with a base learning rate of $3\text{\times}{10}^{-5}$ and an exponential learning rate decay after a 100-step warm-up. The batch size is set to 2, with gradient accumulation over 16 steps for an effective batch size of 32. This is a deliberate strategy to allow for mixing of images with different aspect ratios and resolutions. We use a specific mix of indoor and outdoor scenes from both Hypersim [40] (90%) and Virtual KITTI 2 [6] (10%), which was beneficial to the model’s performance. Fine-tuning takes approximately 3 days on a single Nvidia H100 GPU.
@@
Figure 6: Qualitative results of fixed single-step/ensembled multi-step Marigold, and deterministic end-to-end fine-tuned models based on Marigold and GeoWizard. Multi-step results show noise-like artifacts. The predictions of the end-to-end fine-tuned models tend to be more sharp and accurate. Additional qualitative results can be found in Appendix D.
@@ 6 Experimental Evaluation
Comparison with Marigold.
As is evident from Fig. 5, the fixed DDIM scheduler reveals that Marigold’s [26] multi-step denoising is not actually working: instead of improving the depth map with more denoising steps, the performance actually gets worse. This is because repeatedly denoising sharpens the depth map, but also accumulates errors because the model expects noised ground truth latents instead of its own predictions. This behavior was previously masked by the large error due to the broken DDIM scheduler. Note, however, that the fixed model performs strictly better than before, for any given number of steps. Ensembling does still provide noticeable benefits when using at least two inference steps. For single-step inference, the predictions tend to be highly correlated, in which case ensembling does not lead to significant improvements.
We compare vanilla Marigold and a variant distilled into a Latent Consistency Model (LCM) [34] for few-step inference with our single-step variants in Tab. 1. Using 4 steps and ensemble size 5, Marigold LCM performs on par with the best setting of vanilla Marigold (50 steps, 10 ensemble) on outdoor data (KITTI [17], DIODE [48]), but slightly worse for the indoor datasets. Using a single step only, performance drops even more strongly to 6.5 AbsRel on NYUv2 [46], compared to the previous best 5.5 AbsRel at 50 steps with an ensemble of 10 predictions. In contrast, we see that vanilla Marigold with the fixed DDIM scheduler already reaches 5.7 AbsRel in a single step and without ensembling. Notably, this is already better than the 6.0 AbsRel of 50-step Marigold with the same model weights, and is already very close to Marigold’s best setting with ensembling. Moreover, we show that further end-to-end fine-tuning of Marigold leads to a substantial improvement of -0.5 AbsRel on NYUv2, and surpassing any previous setting of Marigold, all in a single step and without ensembling. Finally, directly fine-tuning Stable Diffusion [41] instead of the Marigold-pretrained model leads to comparable results. Tab. 2 paints the same overall picture for surface normal estimation; the fixed single-step Marigold outperforms the vanilla multi-step, ensembled Marigold, and end-to-end fine-tuning yields even better results, even when starting from Stable Diffusion directly.
@@
Deterministic or Probabilistic.
We perform an ablation on the type of noise used during fixed-timestep fine-tuning; the results are shown in Tab. 4. “Gaussian” and “Pyramid” refer to the standard normal and multi-resolution noise commonly employed in diffusion training and used in Marigold, respectively. “Zeros” describes our default setting, i.e., no noise. We find that using constant zeros performs slightly better than the alternatives, although the method seems to be fairly robust to the actual choice of noise.
Comparison with GeoWizard.
GeoWizard [16] jointly predicts depth and surface normals, thus we also jointly fine-tune the model end-to-end for both tasks. Tab. 3 shows substantial improvements for surface normal estimation. In particular, the fine-tuned model performs substantially better than both the fixed single-step model and the claimed previous best results with 50 steps and ensembling of 10 predictions, which we were not able to reproduce. For depth, we observe smaller, but consistent improvements. We provide more results in Appendix C.
Qualitative Results.
Fig. 6 shows several qualitative examples. The single-step model fails to produce sharp results, while the multi-step ensemble method starts to hallucinate high frequency details. The end-to-end fine-tuned models predict sharp depth maps and high-quality normals.
State-of-the-art Landscape.
As shown in Sec. 6, the fine-tuned models outperform current state-of-the-art generative methods for depth estimation on most datasets. Among discriminative methods, only Depth Anything [51, 52] and Metric3D [54, 23] demonstrate superior performance; however, these methods were trained on datasets that are two to three orders of magnitude larger. For surface normal estimation, the fine-tuned models set new state-of-the-art results across all evaluated datasets, with the exception of NYUv2, where Metric3D v2 continues to lead, as shown in Sec. 6.
@@ 7 Conclusion
Acknowledgements. Karim Knaebel’s research is funded by the Bosch-RWTH LHC project “Context Understanding for Autonomous Systems”. Christian Schmidt was funded by BMBF project bridgingAI (16DHBKI023). Computations were performed with computing resources granted by RWTH Aachen University under project rwth1690.
Supplemental Material
Appendix A Implementation and Evaluation Details
Table 7: Comparison of leading vs. trailing timestep selection. The timesteps selected by two DDIM scheduler timestep selection strategies for $T=1000$ timesteps and varying numbers of inference steps.

### v2

@@ 2 Related Work
Monocular Depth Estimation.  Monocular depth estimation models predict a pixel-wise depth map of a scene from a single image. The most comprehensive representation is metric depth, which requires modeling the focal length to account for different cameras, introducing additional uncertainty [4, 34, 53, 21].
An alternative to metric depth is affine-invariant depth, which is equivalent to metric depth up to an unknown global scale and shift of the scene. This is the representation of choice for a wide range of monocular depth estimation methods [55, 37, 36, 10, 22, 50, 51]. Unlike metric depth, affine-invariant depth is independent of the focal length and is easier to regress in unfamiliar scenes, where no object can serve as a metric reference. However, it still preserves the distance ratios between objects. Metric depth can be recovered from affine-invariant depth by anchoring it with sparse, known depth values or by explicitly estimating the missing scale and shift [54].
@@
The transition from CNNs to ViTs [8] in DPT [36] and Omnidata [10] marked another key advancement in the field. Recent methods such as Depth Anything [50, 51] and Metric3D [53, 21] have followed the success of MiDaS by utilizing the high-capacity ViT-g DINOv2 [33] backbone and training on vast datasets—62M and 16M samples, respectively—one to two orders of magnitude larger than previous work. Notably, Depth Anything retains a simple DPT architecture, but combining a DINOv2 initialization and a large training dataset enables it to generalize to in-the-wild scenarios. Metric3D additionally utilizes the focal length to boost performance, which however limits its training data to scenes with known focal length.
Monocular Normal Estimation.  Surface normal estimation involves predicting the orientation of surfaces in a scene from an image, resulting in a 3D vector representing the surface’s orientation for each pixel. Hoiem et al. [19, 20] were among the pioneers to introduce learning-based approaches for surface normal estimation. Since then, deep learning methods have become dominant, with many notable contributions [47, 11, 3, 1, 10, 22, 2]. Among these, Omnidata [10] stands out for training a UNet-based model on a large-scale dataset of 12M images captured in diverse environments. Its successor, Omnidata v2 [22], advanced the field by transitioning to a ViT-based architecture with a DPT head, incorporating sophisticated 3D-aware data augmentations to enhance generalization. In contrast, DSINE [2] adopts a more data-efficient approach that focuses on introducing new inductive biases to enhance performance. Lastly, Metric3D v2 [21], mentioned earlier, is also capable of predicting surface normals in addition to depth.
Diffusion Models for Geometry Estimation.  Several recent generative text-to-3D methods [30, 35] explicitly produce multi-view depth and normal maps. However, these methods focus on isolated single-object scenarios, making them unsuitable for complex, in-the-wild environments.
Other approaches have utilized diffusion models for scene-level depth estimation [42, 41, 9, 56]. Among these, VPD [56] leverages Stable Diffusion [39] as both an image and a text feature extractor, incorporating a depth regression head that utilizes multi-scale image feature maps alongside text-to-image cross-attention maps. However, these models have not demonstrated robust generalization.
@@
In our work, we observe that Marigold and follow-up methods, aside from DepthFM and Marigold LCM, which are designed for few-step prediction, suffer from a flawed implementation [28] of the DDIM [45] inference pipeline that prevents them from functioning effectively in the few-step regime. Furthermore, although the denoising diffusion fine-tuning objective used by Marigold and follow-up works for depth and normals estimation has shown effectiveness, we find it to be neither a key factor for good results nor clearly superior to task-specific end-to-end fine-tuning.
@@ 3.2 Marigold
Marigold casts depth estimation as a conditional latent diffusion process, allowing it to build upon large pretrained diffusion models such as Stable Diffusion [24]. Marigold is conditioned on images, so following the above argument, we expect its single-step prediction to be a sensible, but blurry depth map.
Training.  Fig. 3 shows an overview of Marigold’s training procedure. Marigold adapts the SD v2 [39] UNet architecture for conditional depth map generation, using $\bm{\mathbf{v}}$-parametrization [40] during training. The training objective is formulated in the latent space of the frozen SD VAE. The GT depth map $\bm{\mathbf{d}}^{*}$ is replicated along the channel dimension to conform to the 3-channel inputs of the VAE and encoded as $\bm{\mathbf{z}}^{*}=\mathcal{E}(\bm{\mathbf{d}}^{*})$. Similarly, the RGB latent is $\bm{\mathbf{x}}=\mathcal{E}(\bm{\mathbf{I}}_{\text{RGB}})$. During training, noise is added only to the depth latent to get $\bm{\mathbf{z}}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{\mathbf{z}}^{*}+\sqrt{1-\bar{\alpha}_{t}}\bm{\mathbf{\epsilon}}$. The UNet receives the concatenated latents and the timestep $t$ as inputs and predicts $\bm{\mathbf{\hat{v}}}_{t}=\bm{\mathbf{\hat{v}}}_{\theta}([\bm{\mathbf{z}}_{t},\bm{\mathbf{x}}],t)$. The first convolutional layer is duplicated to accomodate the larger number of input channels [24].
@@
Fixing Single-Step Inference.  We now turn to analyze the behavior of Marigold during inference. First, we show the pixel-wise standard deviation across steps for Marigold’s default 50-step inference in  Fig. 2. We observe very small differences for almost all pixels, indicating that the model changes predictions very little during inference; in particular, this means that the first prediction of the 50-step schedule already corresponds closely to the final output. While we would expect that the single-step output should be similar to the first step of the 50-step schedule, as shown in Fig. 2 it rather corresponds to pure noise.
@@
Fixing the flaw is simple: we need to align the timestep with the noise level. To do this, we can use the trailing setting as proposed in recent work [28] for image-generative models. However, we emphasize that while this setting provided only slight improvements for image generation [28], it is crucial for single-step inference in models such as Marigold. In Fig. 2 and Fig. 2, we compare the outputs of the same model, using the flawed and the fixed inference schedule, respectively. Clearly, the fixed inference process produces sensible predictions, while the original does not.
@@ 5 Experimental Setup
Training Datasets.  To allow for direct comparison with Marigold [24], we use the same training datasets: Hypersim [38], which consists of photorealistic indoor scenes, and Virtual KITTI 2 [6], which covers driving scenarios. Both datasets are fully synthetic and provide high-quality ground-truth annotations.
@@
Evaluation Datasets.  We evaluate the fine-tuned models on commonly used benchmarks for monocular depth estimation. NYUv2 [44] and ScanNet [7] provide RGB-D data of indoor environments captured with Kinect cameras. ETH3D [43] and DIODE [46] consist of both indoor and outdoor scenes, derived from LiDAR sensors. KITTI [15] contains outdoor driving scenes captured by vehicle-mounted cameras and LiDAR sensors. For surface normal estimation, we evaluate on NYUv2, ScanNet, and additionally on iBims-1 [26], a high-quality indoor RGB-D dataset, as well as Sintel [5], a synthetic outdoor dataset.
Evaluation Protocol.  All evaluations are conducted in the zero-shot setting. We evaluate affine-invariant depth predictions using the standard approach, which involves the same scale and shift optimization between the predicted depth and the ground truth as in the loss computation [37]. We report the mean absolute relative error (AbsRel), defined as the average relative difference between the ground-truth depth and the aligned predicted depth at each pixel, as well as the $\delta$1 accuracy, which is the percentage of pixels where the ratio of the aligned predicted depth to the ground truth (and its inverse) is less than 1.25.
@@
Implementation Details.  For depth estimation, we use the official Marigold checkpoint, whereas for normal estimation, we train a model with the same training setup as Marigold’s depth estimation, encoding normal maps as 3D vectors in the color channels.
Unless noted otherwise, we follow Marigold’s hyperparameters. We train all models for 20K iterations using the AdamW optimizer [31] with a base learning rate of $3\text{\times}{10}^{-5}$ and an exponential learning rate decay after a 100-step warm-up. The batch size is set to 2, with gradient accumulation over 16 steps for an effective batch size of 32. This is a deliberate strategy to allow for mixing of images with different aspect ratios and resolutions. We use a specific mix of indoor and outdoor scenes from both Hypersim [38] (90%) and Virtual KITTI 2 [6] (10%), which was beneficial to the model’s performance. Fine-tuning takes approximately 3 days on a single Nvidia H100 GPU.
@@
Figure 6: Qualitative results of fixed single-step/ensembled multi-step Marigold, and deterministic end-to-end fine-tuned models based on Marigold and GeoWizard. Multi-step results show noise-like artifacts. The predictions of the end-to-end fine-tuned models tend to be more sharp and accurate. Additional qualitative results can be found in the supplementary materials.
@@ 6 Experimental Evaluation
Comparison with Marigold.  As is evident from Fig. 5, the fixed DDIM scheduler reveals that Marigold’s [24] multi-step denoising is not actually working: instead of improving the depth map with more denoising steps, the performance actually gets worse. This is because repeatedly denoising sharpens the depth map, but also accumulates errors because the model expects noised ground truth latents instead of its own predictions. This behavior was previously masked by the large error due to the broken DDIM scheduler. Note, however, that the fixed model performs strictly better than before, for any given number of steps. Ensembling does still provide noticeable benefits when using at least two inference steps. For single-step inference, the predictions tend to be highly correlated, in which case ensembling does not lead to significant improvements.
We compare vanilla Marigold and a variant distilled into a Latent Consistency Model (LCM) [32] for few-step inference with our single-step variants in Tab. 1. Using 4 steps and ensemble size 5, Marigold LCM performs on par with the best setting of vanilla Marigold (50 steps, 10 ensemble) on outdoor data (KITTI [15], DIODE [46]), but slightly worse for the indoor datasets. Using a single step only, performance drops even more strongly to 6.5 AbsRel on NYUv2 [44], compared to the previous best 5.5 AbsRel at 50 steps with an ensemble of 10 predictions. In contrast, we see that vanilla Marigold with the fixed DDIM scheduler reaches 5.7 AbsRel in a single step and without ensembling. Notably, this is better than the 6.0 AbsRel of 50-step Marigold with the same model weights. Moreover, we show that further end-to-end fine-tuning of Marigold leads to a substantial improvement of -0.5 AbsRel on NYUv2, surpassing all previous settings of Marigold, in a single step and without ensembling. Finally, directly fine-tuning Stable Diffusion [39] instead of the Marigold-pretrained model leads to comparable results. Tab. 2 paints the same overall picture for surface normal estimation; fixed single-step Marigold outperforms the vanilla multi-step, ensembled Marigold, and end-to-end fine-tuning yields even better results, even when applied to Stable Diffusion directly.
@@
Deterministic or Probabilistic.  We perform an ablation on the type of noise used during fixed-timestep fine-tuning; the results are shown in Tab. 4. “Gaussian” and “Pyramid” refer to the standard normal and multi-resolution noise commonly employed in diffusion training and used in Marigold, respectively. “Zeros” describes our default setting, i.e., no noise. We find that using constant zeros performs slightly better than the alternatives, although the method seems to be fairly robust to the actual choice of noise.
Comparison with GeoWizard.  GeoWizard [14] jointly predicts depth and surface normals, thus we also jointly fine-tune the model end-to-end for both tasks. Tab. 3 shows substantial improvements for surface normal estimation. In particular, the fine-tuned model performs substantially better than both the fixed single-step model and the claimed previous best results with 50 steps and ensembling of 10 predictions, which we were not able to reproduce. For depth, we observe smaller, but consistent improvements. We provide more results in the supplementary materials.
Qualitative Results.  Fig. 6 shows several qualitative examples. The single-step model fails to produce sharp results, while the multi-step ensemble method starts to hallucinate high frequency details. The end-to-end fine-tuned models predict sharp depth maps and high-quality normals.
State-of-the-art Landscape.  As shown in Sec. 6, the fine-tuned models outperform current state-of-the-art generative methods for depth estimation on most datasets. Among discriminative methods, only Depth Anything [50, 51] and Metric3D [53, 21] demonstrate superior performance; however, these methods were trained on datasets that are two to three orders of magnitude larger. For surface normal estimation, the fine-tuned models set new state-of-the-art results across all evaluated datasets, with the exception of NYUv2, where Metric3D v2 continues to lead, as shown in Sec. 6.
@@ 7 Conclusion
Acknowledgements. Karim Knaebel’s research is funded by the Bosch-RWTH LHC project “Context Understanding for Autonomous Systems”. Christian Schmidt is funded by BMBF project bridgingAI (16DHBKI023). Computations were performed with computing resources granted by RWTH Aachen University under project rwth1690.
Supplementary Material
Appendix A DDIM Inference
During training, the highest noise level corresponds to the last timestep $t=T$, and $t=1$ corresponds to a very small noise level. The DDIM inference scheduler iterates over a series of $k$ timesteps $\tau_{1}>\tau_{2}>\ldots>\tau_{k}>0$ and iteratively denoises the initial noise input $\bm{\mathbf{z}}_{\tau_{1}}$. We consider the leading and trailing schedules that are also discussed by Lin et al. [28] and show the selected timesteps for different $k$ in LABEL:tab:leading_vs_trailing. The original leading timestep selection strategy of the DDIM scheduler excludes the final timestep $T$. This leads to a mismatch between training and inference; using the leading schedule, the model receives noise as input, even though the timestep embedding indicates a partially denoised input. In contrast, the fixed trailing strategy always starts with $t=T$ for the first denoising step, properly aligning training and inference. In the limit of $k\rightarrow T$ inference steps, both strategies converge to the same behavior.
Table A-1: Comparison of leading vs. trailing timestep selection. The timesteps selected by two DDIM scheduler timestep selection strategies for $T=1000$ timesteps and varying numbers of inference steps.